OpenStudy (anonymous):

What is the y-value for f(x) = tan(x) when x = -45°?

OpenStudy (astrophysics):

\[-45 \deg = -\frac{ \pi }{ 4 }\] right?

OpenStudy (astrophysics):

\[\tan \theta = \frac{ \sin \theta }{ \cos \theta }\]

OpenStudy (anonymous):

I think so thank you!

OpenStudy (astrophysics):

 use this triangle to solve your question